During 2019, the Tastee Partnership reported income before guaranteed payments of $92,000. Stella owns a 90% profits interest and works 1,600 hours per year in the business. Euclid owns a 10% profits interest and performs no services for the partnership during the year. For services performed in 2019, Stella receives a "salary" of $6,000 per month. Euclid withdrew $10,000 from the partnership during the year. a. What is the amount of guaranteed payments made by the partnership during 2019? b. How much is the partnership's ordinary income after any deduction for guaranteed payments? c. For 2019, how much income will Stella and Euclid report?
Answer #1

a) Amount of guaranteed payments made by partnership = Salary paid for services performed by Stella

= 12 x 6000

= $ 72000

b) Partnership's ordinary income after deduction of guaranteed payment= 92000 - 72000 =$ 20,000

c) Stella own = 90% of 20000 = 18000 and Euclid own = 10 % of 20000 = 2000.Since Euclid withdrew already 10000 and from partnership in excess of her share limit.

Difference of 10000 - 2000 = 8000 is distributed as per agrement which is 90% for Stella

Stella income Report = 90% of 20,000 + Her salary = 18000 + 72000 - ( 90% of 8000) = 82800.

Euclid income Report = $ 10000
Answer #2

Euclid income reported is $2,000 not $10,000. The $10,000 he withdrew from the year is not deductable by partnership nor taxable to euclid.

source: tax law
answered by: a